OpenStudy (anonymous):

In a shipment of 400 parts, 14 are found to be defective. How many defective parts should be expected in a shipment of 1000?

OpenStudy (anonymous):

multiply 14 by 2.5 because 2.5 400s go into 1000

OpenStudy (anonymous):

you get 35 as an answer

OpenStudy (anonymous):

okay 400=14 so 800=14+14=28 you still have 200 left over which is half of 400 so you can take 14/2=7 so you know 14+14+7=35